# DyRoNet: A Low-Rank Adapter Enhanced Dynamic Routing Network for Streaming Perception

### 15 Mar 2024 v2

@@ Abstract
The quest for real-time, accurate environmental perception is pivotal in the evolution of autonomous driving technologies. In response to this challenge, we present DyRoNet, a Dynamic Router Network that innovates by incorporating low-rank dynamic routing to enhance streaming perception. DyRoNet distinguishes itself by seamlessly integrating a diverse array of specialized pre-trained branch networks, each meticulously fine-tuned for specific environmental contingencies, thus facilitating an optimal balance between response latency and detection precision. Central to DyRoNet’s architecture is the Speed Router module, which employs an intelligent routing mechanism to dynamically allocate input data to the most suitable branch network, thereby ensuring enhanced performance adaptability in real-time scenarios. Through comprehensive evaluations, DyRoNet demonstrates superior adaptability and improved performance over existing methods, efficiently catering to a wide variety of environmental conditions and setting new benchmarks in streaming perception accuracy and efficiency. Beyond establishing a paradigm in autonomous driving perception, DyRoNet also offers engineering insights and lays a foundational framework for future advancements in streaming perception. For further information and updates on the project, visit https://tastevision.github.io/DyRoNet/.
@@ 1 Introduction
In autonomous driving, achieving low-latency and high-precision perception stands as a critical challenge. Traditional object detection algorithms Zou et al. (2023), though effective across a variety of scenarios, often face latency issues due to high computational demands. This latency leads to a mismatch between algorithmic predictions and the real-world positions of objects, adversely affecting detection accuracy and the decision-making process crucial for safe autonomous navigation Chen et al. (2023). To address these challenges, streaming perception has emerged as a groundbreaking approach, offering a novel perspective on object detection Li et al. (2020). Diverging from traditional methods that concentrate on immediate detection, streaming perception endeavors to predict future environmental conditions, thus aiming to reduce the lag between computation and real-world dynamics. This approach enhances the real-time responsiveness of autonomous driving systems, marking a leap towards adaptive, efficient, and safer vehicular autonomy Hauptmann et al. (2023); Muhammad et al. (2020).
@@
Despite its advantages, streaming perception’s application in unpredictable and dynamic autonomous driving reveals key challenges. The performance of perception systems is laboriously influenced by factors such as camera motion, weather variability, lighting changes Zhu et al. (2024), and the presence of small objects. These factors introduce performance inconsistencies, challenging the robustness and reliability of a singular, one-size-fits-all model approach (see Sec. 3.1). Such diversity in driving environments mandates the development of more adaptable and versatile perception models Guo et al. (2019):
(1) Adaptation to Diverse Scenarios: The wide range of conditions encountered in autonomous driving requires models capable of adjusting to an extensive variety of scenarios (refer to Fig. 1). Crafting algorithms that maintain their core effectiveness while being adaptable to specific environmental conditions presents a significant challenge, further discussed in Sec. 3.1 and Appendix A.
(2) Balancing Accuracy and Efficiency: Achieving a balance between precision and processing speed when employing models of varying complexities is vital. High-detail models may enhance accuracy but at the cost of increased latency, while simpler models might offer faster inference times but reduced precision. Identifying the optimal compromise between these extremes is essential for enhancing streaming perception’s effectiveness, as Sec. 3.1 and Sec. 4.3.
Addressing the critical challenges of diverse scenario distribution and performance-efficiency balance is paramount for the progression of autonomous driving technologies, a theme extensively discussed in Sec. 3.1. In response, we introduce DyRoNet, a tailored framework designed to enhance the dynamic routing capabilities essential for autonomous systems. This framework, developed specifically for streaming perception, integrates a low-rank dynamic routing mechanism enriched with a sophisticated speed router module. This central module adeptly directs incoming data to the most suitable pre-trained branch network, with each branch meticulously optimized for peak performance across a spectrum of environmental conditions. Detailed operational insights and the strategic underpinnings of DyRoNet are elaborated upon in Sec. 3.2, highlighting its key contribution to augmenting the adaptability and efficiency of autonomous driving systems. The main contributions can be summarized as:
@@
  • Environmental Velocity Impact: Our research emphasizes the key role of environmental speed in streaming perception. Through an in-depth analysis of various environmental factors, we demonstrate the essential need for adaptive perception systems. This insight significantly advances the understanding of how dynamic environmental conditions influence streaming perception in autonomous driving.
  • Innovation of Speed Router: At the core of DyRoNet is the innovative ‘Speed Router’, a specially designed component that dynamically identifies the most suitable processing route for incoming data. This feature not only enhances the efficiency and accuracy of perception tasks but also exhibits exceptional adaptability to changing environmental conditions.
  • Comprehensive Evaluation: Through extensive testing of DyRoNet using diverse branch selection strategies, we highlight its unmatched adaptability and improved performance across different scenarios. This extensive evaluation confirms the framework’s effectiveness and its capacity to set new benchmarks in low-latency, high-accuracy streaming perception.
@@
In summary, DyRoNet thus marks a paradigm shift in the pursuit of low-latency, high-accuracy perception for autonomous driving, directly addressing the foremost challenges of environmental adaptability and dynamic branch processing in streaming perception.

### 18 Mar 2024 v3

# DyRoNet: Dynamic Routing and Low-Rank Adapters for Autonomous Driving Streaming Perception

@@ Abstract
The advancement of autonomous driving systems hinges on the ability to achieve low-latency and high-accuracy perception. To address this critical need, this paper introduces Dynamic Routering Network (DyRoNet), a low-rank enhanced dynamic routing framework designed for streaming perception in autonomous driving systems. DyRoNet integrates a suite of pre-trained branch networks, each meticulously fine-tuned to function under distinct environmental conditions. At its core, the framework offers a speed router module, developed to assess and route input data to the most suitable branch for processing. This approach not only addresses the inherent limitations of conventional models in adapting to diverse driving conditions but also ensures the balance between performance and efficiency. Extensive experimental evaluations demonstrating the adaptability of DyRoNet to diverse branch selection strategies, resulting in significant performance enhancements across different scenarios. This work not only establishes a new benchmark for streaming perception but also provides valuable engineering insights for future work. [^1] Project: https://tastevision.github.io/DyRoNet/
@@ 1 Introduction
In autonomous driving systems, it is crucial to achieve low-latency and high-precision perception. Traditional object detection algorithms Zou et al. (2023), while effective in various contexts, often confront the challenge of latency due to inherent computational delays. This lag between algorithmic processing and real-world states can lead to notable discrepancies between predicted and actual object locations. Such latency issues have been extensively reported and are known to significantly impact the decision-making process in autonomous driving systems Chen et al. (2023).
Addressing these challenges, the concept of streaming perception has been introduced as a response Li et al. (2020). This perception task aims to predict “future” results by accounting for the delays incurred during the frame processing stage. Unlike traditional methods that primarily focus on detection at a given moment, streaming perception transcends this limitation by anticipating future environmental states, and aligning perceptual outputs closer to real-time dynamics. This new paradigm is key in addressing the critical gap between real-time processing and real-world changes, thereby enhancing the safety and reliability of autonomous driving systems Muhammad et al. (2020).
@@
Although the existing streaming approach seems promising, it still faces contradictions in real-world scenarios. These contradictions primarily stem from the diverse and unpredictable nature of driving environments. The factors such as camera motion, weather conditions, lighting variations, and the presence of small objects seriously impact the performance of perception measures, leading to fluctuations that challenge their robustness and reliability (see Sec. 3.1). This complexity in real-world scenarios underscores the limitations of a single, uniform model, which often struggles to adapt to the varied demands of different driving conditions Guo et al. (2019). In general, the challenges of streaming perception mainly include:
(1) Diverse Scenario Distribution: Autonomous driving environments are inherently complex and dynamic, showing a myriad of scenarios that a single perception model may not adequately address (see Fig. 1). The need to customize perception algorithms to specific environmental conditions, while ensuring that these models operate cohesively, poses a significant challenge. As discussed in Sec. 3.1, adapting models to various scenarios without compromising their core functionality is a crucial aspect of streaming perception.
(2) Performance-Efficiency Balance: To our knowledge, the integration of both large and small-scale models is essential to handle the varying complexities encountered in different driving scenes. The large models, while potentially more accurate, may suffer from increased latency, whereas smaller models may offer faster inference at the cost of reduced accuracy. Balancing performance and efficiency, therefore, becomes a challenging task. In Sec. 3.1, we explore the strategies for optimizing this balance, exploring how different model architectures can be effectively utilized to enhance streaming perception.
Generally speaking, these challenges highlight the demand for streaming perception. As we study in Sec. 3.1, addressing the diverse scenario distribution and achieving an optimal balance between performance and efficiency are key to advancing the state-of-the-art in autonomous driving. To address the intricate challenges presented by real-world streaming perception, we introduce DyRoNet, a framework designed to enhance dynamic routing capabilities in autonomous driving systems. DyRoNet stands as a low-rank enhanced dynamic routing framework, specifically crafted to cater to the requirements of streaming perception. It encapsulates a suite of pre-trained branch networks, each meticulously fine-tuned to optimally function under distinct environmental conditions. A key component of DyRoNet is the speed router module, ingeniously developed to assess and efficiently route input data to the most appropriate branch, as detailed in Sec. 3.2. To sum up, the contributions are listed as:
@@
  • We emphasize the impact of environmental speed as a key determinant of streaming perception. Through analysis of various environmental factors, our research highlights the imperative need for adaptive perception responsive to dynamic conditions.
  • By utilizing a variety of sophisticated streaming perception techniques, DyRoNet provides the speed router as a major invention. This component dynamically determines the best route for handling each input, ensuring efficiency and accuracy in perception. The ability to adapt and be versatile is demonstrated by this dynamic route-choosing mechanism.
  • Extensive experimental evaluations have demonstrated that DyRoNet is capable of adapting to diverse branch selection strategies, resulting in a substantial enhancement of performance across various branch structures. This not only validates the framework’s wide-ranging applicability but also confirms its effectiveness in handling different real-world scenarios.
@@
In summary, DyRoNet offers advancements for low-latency, high-accuracy perception in autonomous driving. By addressing challenges of environmental adaptability and dynamic branch selection, DyRoNet sets new benchmarks in achieving low-latency and high-accuracy perception.